OpenStudy (anonymous):

Solve z^3 = 4 sqrt(2*(i-1)). I get nowhere.

OpenStudy (anonymous):

Please, is there no one who fancy some quizzy math prob?

OpenStudy (anonymous):

@mathmale , could you maybe take a look?

OpenStudy (anonymous):

@Lena772 ?

OpenStudy (lena772):

@thomaster

OpenStudy (anonymous):

Ummm, consider \((a+bi)^3=4\sqrt{2(i-1)}\)

OpenStudy (anonymous):

I'd start by squaring both sides: \[ (a+bi)^3=16(2(i-1)) = -32+32i \]

OpenStudy (anonymous):

\[ (a+bi)^2 =a^2+2abi - b^2 \]

OpenStudy (anonymous):

\[ (a+bi)(a^2-b^2+2abi) = (\ldots) +(\ldots)i \]

OpenStudy (ranga):

When left side is squared it will be (a + bi)^6

hartnn (hartnn):

how did u stumble upon this ugly question ?

hartnn (hartnn):

i would try to break this problem in 2 parts first find x+iy = 4 sqrt(2*(i-1)) square both sides and compare real and imaginary parts to get x and y then solving z^3 = x+iy

OpenStudy (ranga):

z^6 = 32(-1 + i) z = { 32(-1 + i) }^1/6 = 2^5/6 * (-1 + i)^1/6 Put (-1 + i) in r(cos(theta) + isin(theta)) Then use DeMoivre's Theorem to raise it to ^1/6.

OpenStudy (dumbcow):

\[-1 + i = \sqrt{2}(\cos \frac{3\pi}{4} + i \sin \frac{3\pi}{4})\]

OpenStudy (anonymous):

 now solve.